# Language Guided Skill Discovery

### v1

@@ 4.3 Language distance as a valid metric
We begin with defining $d_{\text{lang}}$ formally. Let us denote a user prompt as $l_{\text{prompt}}$, and a rule-based annotator as $g:\mathcal{S}\to l_{\text{state}}$. The annotator $g$ is a pre-defined function that translates state vector into corresponding natural language sentence without any addition or loss of information. Then, we denote the output description of an LLM for state $s$ as $l_{\text{desc}}(s)=LLM(\cdot|l_{\text{prompt}},l_{\text{state}})$. We configure the LLM to be deterministic with zero temperature during generation, thus treating it as a function. Next, we employ a pre-trained language embedding model $f_{\text{embd}}:l_{\text{desc}}\to\mathbb{R}^{N}$. Using this setup, we can now define the language distance $d_{\text{lang}}$ as follows:
@@
We provide a proof of the claim in Appendix B. Applying this modification, we came to the final objective function as follows:
@@ 4.4 Utilizing learned skills using natural language
After training is complete, we need a way of selecting specific $z$ values to utilize the behaviors associated with each skill to solve concrete tasks. However, we lack knowledge about which $z$ leads to the desired behavior. We can sweep over the skill space at fixed intervals [Laskin et al., 2022], or sample a number of skills and evaluate each with hundreds of rollout trajectories [Park et al., 2023a]. To overcome these challenges, Park et al. [2021, 2023b] have suggested using $z=(\phi(g)-\phi(s))/|\phi(g)-\phi(s)|_{2}$ for continuous skills, when we have Lipschitz continuous $\phi$. Yet, in many cases, we cannot fully specify the desired goal state. For instance, consider an agent using a robotic arm to move a cube. The state includes the end-effector’s rotation and position, as well as the cube’s coordinates. While you can specify the cube’s position, the required rotation and position of the end-effector to achieve that position may not be known.
To address this challenge of goal state specification, we propose training a separate network $\psi:f_{\text{embd}}(l_{\text{desc}})\to Z$ to infer the skill. Intuitively, $\psi$ is a neural network designed to infer which $z$ was used to reach the state that yields the given description. If we train $\psi$ alongside the goal-conditioned policy $\pi$, then we can use $\psi$ to select $z$, which is then fed into $\pi$. In this way, the agent can reach the desired state as described in natural language in a zero-shot manner. The full algorithm, including the training of $\psi$, is presented in Appendix D.
@@ 5 Experiments
In this section, we evaluate our proposed LGSD by conducting a series of experiments on continuous control environments, encompassing both locomotion and manipulation setups. We aim to answer three questions: (1) Can prompting constrain the skill space into a desired semantic subspace? (2) Can language guidance lead to obtaining more diverse skills compared to unsupervised skill discovery baselines? (3) Can we employ learned skills using natural language?
@@
We trained our algorithm and baselines using Isaac Gym [Makoviychuk et al., 2021], a high-throughput GPU-based physics simulator. For the language model, we employed gpt-4-turbo-2024-04-09[Achiam et al., 2023]. We set the temperature parameter of the language model to 0 to get a consistent, low-variance measure of $d_{\text{lang}}$. To reduce the number of unique queries, we discretized and cached the input and output of these queries and reused them during training. We provide the exact prompts used for each experiments in Appendix F.
@@ 5.1 Constraining the skill space into a desired semantic subspace
We first explored how LGSD constrains the skill space into a desired semantic subspace using prompts. To visually present this concept, we designed experiments where agents visit mutually exclusive states depending on the prompts. We tested this idea in two environments, Ant and FrankaCube. In the Ant environment, we required the agent to traverse only half of the plane. For instance, we allowed an agent to visit any location in the northern part of the plane but restricted it from entering the southern area. We repeated these experiments for each cardinal direction: North, South, East, and West (NSEW). Similarly, in the FrankaCube environment, we intended for the robot arm to move the object to only half of the plane for each of the NSEW directions starting from the origin. We trained four agents, each with its own unique prompt, encouraging the agent to explore only within the intended semantic subspace.
@@
Fig. 4 (and Fig. 9 in Appendix) shows the discovered skills trained with each prompt. For each plot, we sampled $z\sim\mathcal{N}(0,\mathbb{I})$ 150 times independently and then generated trajectories using a skill-conditioned policy. As shown in Fig. 4, LGSD effectively constrained the skill space into the desired subspace, especially in the Ant environment. For the Franka agent, the robot arm was able to move the object into the intended region, but it could not fully cover all skills within the constrained subspace. We believe this limitation is due to the inherent nature of the exploration problem. In the Ant environment, the dynamics of the agent is invariant to horizontal translation, which allows effective control at any location. On the other hand, moving the cube along a long trajectory often requires changes in pushing strategies with different contacts and force profiles.
@@ 5.2 Language guidance aids discovering diverse skills
To observe the effect of language guidance, we evaluate the diversity measure of the skills learned using LGSD against unsupervised skill discovery baselines. For the baselines, we take two classical mutual information-based skill discovery algorithms: DIAYN [Eysenbach et al., 2018] and DADS [Sharma et al., 2019], as well as three state-of-the-art distance maximization based approaches: LSD [Park et al., 2021], CSD [Park et al., 2023a], and METRA [Park et al., 2023b]. Although METRA was originally designed for pixel-based input tasks, it can also handle vector-based inputs. Therefore, we used vector-based state inputs for all methods to facilitate direct comparison. For all methods, we utilized no human prior knowledge in selecting parts of the state dimension to induce desired behaviors; instead, we passed full observations directly to the skill discovery algorithms. Detailed information about the observations is provided in Appendix E.1.
@@
We present the results in Fig. 6 with two diversity criteria. We trained LGSD and all baseline methods with four different seeds, reporting the mean and standard deviation of the results. Fig. 6 suggests that LGSD discovers more diverse states compared to all baseline algorithms, confirming that language guidance from an LLM helps in discovering both diverse and meaningful skills. Notably, all distance-maximization methods (CSD, METRA, LSD) could manipulate the object while MI-based methods could not. We also present the qualitative results of the learned behaviors by each method in Fig. 7, where LGSD was able to manipulate objects to cover more grids. Additionally, we provide visualizations and an analysis of the agent’s movements in the latent space in Appendix C.
5.3 Utilizing learned skills via language
@@
Lastly, we demonstrate how we can design an agent that can execute learned skills to follow a natural language instruction. As discussed in Section 4.4, our idea is to utilize $\psi(z|f_{\text{embed}}(l_{\text{desc}}))$ to infer a skill to reach the goal state. For instance, our instruction can be a description of desired state,"The object is located at [0.3,0.2]", which does not contain information regarding robot-arm’s position or orientation. Nevertheless, the learned $\psi$ allows an agent to translate the given language instruction into a skill to reach the goal state.
Fig. 8 illustrates scenarios that a robot-arm moves the object to the six different goal positions based on language description. This experiment showcased that the agent can understand the desired goal state with the help of $\psi$, execute the skill accurately, and successfully move the cube to the target positions in a zero-shot manner. We also provide the results of the zero-shot goal-following agent in the locomotion domain in Appendix, Fig. 10.
@@ 6 Conclusion
Appendix A Proof of maximizing Wasserstein Dependency Measure
We begin with the alternative form of Wasserstein distance using Kantorovich-Rubenstein duality following Villani et al. [2009], to make $\mathcal{I_{W}}$ tractable.
@@ Appendix A Proof of maximizing Wasserstein Dependency Measure
where $\mathcal{L}_{S\times Z}$ is a set of all 1-Lipschitz function defined in ${S\times Z}\to\mathbb{R}$ under the distance metric $d$. For the remainder, we follow the proof outlined in Park et al. [2023b], but we provide a simplified version.
We first parameterize $f(S,Z)=\phi(S)^{T}\psi(Z)$ using $\phi:\mathcal{S}\to\mathbb{R}^{D}$ and $\psi:\mathcal{Z}\to\mathbb{R}^{D}$, which are 1-Lipschitz constrained functions. This is safe decomposition because the expressive power of $\phi(S)^{T}\psi(Z)$ is universal. Please refer to Park et al. [2023b] for more detail.
@@
Appendix B Proof of Claim 1
@@
    $\forall x,y\in S_{adj},\ \|\phi(x)-\phi(y)\|\leq d_{\text{lang}}(x,y)\implies\forall x,y\in S,\ \|\phi(x)-\phi(y)\|\leq\tilde{d}(x,y)$
@@
We follow the proof of CSD [Park et al., 2023a] similarly, but it has significant differences which will be covered in later of this section.
@@
Now we are ready to move on to the main proof. We assume that $\forall x,y\in S_{adj},\ \|\phi(x)-\phi(y)\|\leq d_{\text{lang}}(x,y)$. Also, note that for all $x,y\in\mathcal{S}$, both $d_{\text{lang}}(x,x)=0$ and $d_{\text{lang}}(x,y)=d_{\text{lang}}(y,x)$ holds. We denote the optimal path as $p^{*}=(x,s_{1}^{*},s_{2}^{*},...,s_{T-1}^{*},y)\in P(x,y)$ which satisfy $\tilde{d}(x,y)=D_{\text{lang}}(p^{*})$. Then,
@@
The primary distinction is that we focus exclusively on the existing path $p\in P(x,y)$. Due to this specific focus, $\tilde{d}$ no longer serves as a lower bound for $d_{\text{lang}}$. We concentrate on the existing path because our real optimization data originates solely from this path. Consequently, we aim to align our theory more closely with the actual training scenario.
@@
Appendix C Latent space visualization
Figure 11: Each sub-figure depicts a trajectory drawn from a policy conditioned on a randomly sampled skill. The left figure displays the actual paths of both the end-effector (blue) and the object (red) in the real world. The right figure illustrates the corresponding trajectory in latent space, generated using the learned $\phi$ function. A black arrow indicates the skill vector used to produce the trajectory. The numbers indicate transition step in the episode.
As discussed in Section 4.1, LGSD encourages the agent to distribute evenly across the latent space, which is regulated by the language distance, $d_{\text{lang}}$. Fig. 11 demonstrates that the agent’s trajectory in the latent space aligns well with the sampled skill vector. Since the skill vector is sampled isotropically, agent effectively spread out in the latent space. Additionally, it is noteworthy that the distance between two consecutive points in the latent space increases especially when the cube is moved. On the other hand, when the cube is stationary, there is no further extension in the latent space as can be seen in the Fig. 11 (a). This indicates that our language distance $d_{\text{lang}}$ serves as a meaningful proxy for semantic distance.
Appendix D Full algorithm of LGSD
@@
4:        Sample skill $z\sim p(z)$
5:        for $k\leftarrow 1$ to #steps per episode do
@@
8:           Query $LLM(\cdot|s,l_{\text{prompt}})$ to get $l_{\text{desc}}(s)$ and $l_{\text{desc}}(s^{\prime})$
@@
11:           Add $\{s,a,r,s^{\prime},d_{lang}(s,s^{\prime}),l_{\text{desc}}(s),l_{\text{desc}}(s^{\prime}),z\}$ to buffer $\mathcal{D}$
12:        end for
13:     end for
14:     Update $\phi$ to maximize $\mathbb{E}_{(s,z,s^{\prime})\sim\mathcal{D}}\left[(\phi(s^{\prime})-\phi(s))^{T}z+\lambda\cdot\min(\epsilon,d_{\text{lang}}(s,s^{\prime})-\lVert\phi(s)-\phi(s^{\prime})\rVert_{2}^{2})\right]$
15:     Update $\lambda$ to minimize $\mathbb{E}_{(s,z,s^{\prime})\sim\mathcal{D}}\left[\lambda\cdot\min(\epsilon,d_{\text{lang}}(s,s^{\prime})-\lVert\phi(s)-\phi(s^{\prime})\rVert_{2}^{2})\right]$
16:     Update $\pi$ using PPO with reward $r$
17:     Update $\psi$ to minimize Mean_Squared_Error$(\psi(f_{\text{embed}}(l_{\text{desc}}(s))),z)$
18:  end for
Appendix E Implementation details
E.1 Observations

### v2

@@ 4.3 Language distance as a valid metric
We begin with defining $d_{\text{lang}}$ formally. Let us denote a user prompt as $l_{\text{prompt}}$, and a rule-based state annotator as $g:\mathcal{S}\to l_{\text{state}}$. The annotator $g$ is a pre-defined function that translates state vector into corresponding natural language sentence without any addition or loss of information. Then, we denote the output description of an LLM for state $s$ as $l_{\text{desc}}(s)=LLM(\cdot|l_{\text{prompt}},l_{\text{state}})$. We configure the LLM to be deterministic with zero temperature during generation, thus treating it as a function. Next, we employ a pre-trained language embedding model $f_{\text{embd}}:l_{\text{desc}}\to\mathbb{R}^{N}$. Using this setup, we can now define the language distance $d_{\text{lang}}$ as follows:
@@
We provide a proof of the claim in Appendix C. Applying this modification, we came to the final objective function as follows:
@@ 4.4 Utilizing learned skills using natural language
After training is complete, we need a way of selecting specific $z$ values to utilize the behaviors associated with each skill to solve concrete tasks. However, we lack knowledge about which $z$ leads to the desired behavior. We can sweep over the skill space at fixed intervals (Laskin et al., 2022), or sample a number of skills and evaluate each with hundreds of rollout trajectories (Park et al., 2023a). To overcome these challenges, Park et al. (2021; 2023b) have suggested using $z=(\phi(g)-\phi(s))/|\phi(g)-\phi(s)|_{2}$ for continuous skills, when we have Lipschitz continuous $\phi$. Yet, in many cases, we cannot fully specify the desired goal state. For instance, consider an agent using a robotic arm to move a cube. The state includes the end-effector’s rotation and position, as well as the cube’s coordinates. While you can specify the cube’s position, the required rotation and position of the end-effector to achieve that position may not be known.
To address the challenge of goal state specification, we propose training a separate network $\psi:f_{\text{embd}}(l_{\text{desc}}(s))\to Z$ to infer the skill. Intuitively, $\psi$ takes the embedding of the language description of the state $s$ as input and infers which $z$ was used to reach that specific state. Since we are gathering pairs of state $s$ and the corresponding skill $z$ during the training of $\pi$, we can train $\psi$ using the same data. After training, we can use $\psi$ and $s_{\text{goal}}$ to produce $z_{\text{goal}}$, which enables the agent to reach $s_{\text{goal}}$. We then feed $z_{\text{goal}}$ into $\pi$, allowing the agent to reach the goal state as described in natural language in a zero-shot manner. The full algorithm, including the training of $\psi$, is presented in Appendix E.
@@ 5 Experiments
In this section, we evaluate our proposed LGSD by conducting a series of experiments on continuous control environments, encompassing both locomotion and manipulation setups. We aim to answer four questions: (1) Can prompting constrain the skill space into a desired semantic subspace? (2) Can language guidance lead to obtaining more diverse skills compared to unsupervised skill discovery baselines? (3) Can we utilize learned skills for solving downstream tasks? (4) Can we employ learned skills using natural language?
@@
We trained our algorithm and baselines using Isaac Gym (Makoviychuk et al., 2021), a high-throughput GPU-based physics simulator. For the language model, we employed gpt-4-turbo-2024-04-09(Achiam et al., 2023). We set the temperature parameter of the language model to 0 to get a consistent, low-variance measure of $d_{\text{lang}}$. To reduce the number of unique queries, we discretized states and cached the input and output of these queries and reused them during training. We provide the exact prompts used for each experiments in Appendix G.
@@ 5.1 Constraining the skill space into a desired semantic subspace
We first explored how LGSD constrains the skill search space into a desired semantic subspace using prompts. To visually present this concept, we designed experiments where agents visit mutually exclusive states depending on the prompts. We tested this idea in two environments, Ant and FrankaCube. In the Ant environment, we required the agent to traverse only half of the plane. For instance, we allowed an agent to visit any location in the northern part of the plane but restricted it from entering the southern area. We repeated these experiments for each cardinal direction: North, South, East, and West (NSEW). Similarly, in the FrankaCube environment, we intended for the robot arm to move the object to only half of the plane for each of the NSEW directions starting from the origin. We trained four agents, each with its own unique prompt, encouraging the agent to explore only within the intended semantic subspace.
Fig. 4 (and Fig. 10 in Appendix) shows the discovered skills trained with each prompt. For each plot, we sampled $z\sim\mathcal{N}(0,I)$ 150 times independently and then generated trajectories using a skill-conditioned policy. As shown in Fig. 4, LGSD effectively constrained the skill space into the desired subspace, especially in the Ant environment. For the Franka agent, the robot arm was able to move the object into the intended region, but it could not fully cover all skills within the constrained subspace. We believe this limitation is due to the inherent nature of the exploration problem. In the Ant environment, the dynamics of the agent is invariant to horizontal translation, which allows effective control at any location. On the other hand, moving the cube along a long trajectory often requires changes in pushing strategies with different contacts and force profiles.
@@ 5.2 Language guidance aids discovering diverse skills
To observe the effect of language guidance, we evaluate the diversity measure of the skills learned using LGSD against unsupervised skill discovery baselines. For the baselines, we take two classical mutual information-based skill discovery algorithms: DIAYN (Eysenbach et al., 2018) and DADS (Sharma et al., 2019), as well as three state-of-the-art distance maximization based approaches: LSD (Park et al., 2021), CSD (Park et al., 2023a), and METRA (Park et al., 2023b).
Although METRA was originally designed for pixel-based input tasks, it can also handle vector-based inputs. Therefore, we used vector-based state inputs for all methods to facilitate direct comparison. For all methods, we utilized no human prior knowledge in selecting parts of the state dimension to induce desired behaviors; instead, we passed full observations directly to the skill discovery algorithms. Detailed information about the observations is provided in Appendix F.2.1.
@@
We present the results in Fig. 6. As shown in Fig. 6, LGSD discovers more diverse states compared to all baseline algorithms. Skills trained using LGSD moved the object an average of 0.5m, which is five times greater than the value achieved by the second-best algorithm. This result confirms that language guidance from an LLM aids in discovering both diverse and meaningful skills. Notably, all distance-maximization methods (CSD, METRA, LSD) could manipulate the object, whereas MI-based methods could not. Fig. 7 shows qualitative results of the learned behaviors by each method, where LGSD demonstrated the ability to manipulate objects to cover more grid areas. Additionally, visualizations and an analysis of the agent’s movements in latent space are provided in Appendix D.
5.3 Harnessing learned skills
Zero-shot task solving via natural language
@@ 5.3 Harnessing learned skills
We first demonstrate how our agent can solve a downstream task in a zero-shot manner via natural language instruction. As discussed in Section 4.4, we utilize $\psi(z|f_{\text{embed}}(l_{\text{desc}}))$ to infer a skill to reach the goal state. For instance, our instruction can be a description of the desired state, ”The object is located at [0.3,0.2]”, which does not contain information regarding the robot-arm’s position or orientation. The learned $\psi$ allows an agent to translate the given language instruction into an effective skill for reaching the goal state.
Fig. 8 illustrates scenarios that a robot-arm moves the object to the six different goal positions described in natural language. This experiment showcased that the agent can understand the desired goal state with the help of $\psi$, execute the skill accurately, and successfully move the cube to the target positions in a zero-shot manner. We also provide the results of the zero-shot goal-following agent in the locomotion domain in Appendix, Fig. 11.
Training high-level controller with learned skills
Figure 9: The training curve of the high-level controller in locomotion tasks, using the low-level controller trained with LGSD and five state-of-the-art unsupervised skill discovery methods. LGSD is able to achieve better returns and exhibits better sample efficiency.
Although our method enables training a desired set of skills to some extent using tailored prompts, it can be challenging for users to specify exact prompts that produce behaviors for more complex tasks. In such scenarios, our method can act as a pre-training module, which trains a low-level controller $\pi_{low}(a|s,z)$, which can then be leveraged by a high-level controller $\pi_{high}(z|s,g)$ to select the appropriate skills for new downstream task, represented by $g$.
We evaluate the performance of high-level controllers that leverage a pre-trained LGSD low-level controller on two downstream tasks: AntSingleGoal and AntMultiGoal. Details of the downstream tasks are available in Appendix F.1. Fig. 9 shows that by using the set of skills trained with our method, the high-level controller can achieve higher returns and better sample complexity compared to prior skill discovery methods. We hypothesize that these improvements comes from the fact that the skills discovered through LGSD can better focus on more semantically meaningful changes in the $x,y$ positions by leveraging guidance from the LLM, while other methods simply seek changes in all 39 dimensions of the states equally.
@@ 6 Conclusion
Reproducibility Statement
We have made significant efforts to ensure the reproducibility of our work across various aspects.
  • Detailed information about the observations used in our experiments is provided in Appendix F.2.1.
  • A comprehensive pseudo-code of our algorithm is available in Appendix E.
  • A rigorous proof of Claim 4.3 is included in Appendix C.
  • Visualizations of the learned representation space can be found in Appendix D.
  • Details of the high-level tasks training are provided in Appendix F.1.
  • The exact version of the LLM used and all prompts are documented in Appendix G.
Acknowledgements
This research has been funded by the Industrial Technology Innovation Program (P0028404, development of a product-level humanoid mobile robot for medical assistance equipped with bidirectional customizable human-robot interaction, autonomous semantic navigation, and dual-arm complex manipulation capabilities using large-scale artificial intelligence models) of the Ministry of Industry, Trade and Energy of Korea.
Appendix
Appendix A Additional Result Figures
(a) Moving North
(b) Moving East
(c) Moving South
(d) Moving West
Figure 10: LGSD constrains different prompts into different semantic subspace of skills. Each yellow line indicates a trajectory generated using a randomly sampled skill learned with each prompt. Using different prompts, each set of skills is constrained to visit the (a)Northern, (b)Eastern, (c)Southern, and (d)Western areas of the plane.
Figure 11: Ant agents move toward the goal. Red dots indicate the actual path of the agent’s base, while the blue dot represents ths goal. Each goal is specified using natural language.
Appendix B Proof of maximizing Wasserstein Dependency Measure
We begin with the alternative form of Wasserstein distance using Kantorovich-Rubenstein duality following Villani et al. (2009), to make $\mathcal{I_{W}}$ tractable.
@@ Appendix B Proof of maximizing Wasserstein Dependency Measure
where $\mathcal{L}_{S\times Z}$ is a set of all 1-Lipschitz function defined in ${S\times Z}\to\mathbb{R}$ under the distance metric $d$. For the remainder, we follow the proof outlined in Park et al. (2023b), but we provide a simplified version.
We first parameterize $f(S,Z)=\phi(S)^{T}\psi(Z)$ using $\phi:\mathcal{S}\to\mathbb{R}^{D}$ and $\psi:\mathcal{Z}\to\mathbb{R}^{D}$, which are 1-Lipschitz constrained functions. This is safe decomposition because the expressive power of $\phi(S)^{T}\psi(Z)$ is universal. Please refer to Park et al. (2023b) for more detail.
@@
Appendix C Proof of Claim 1
@@
    $\forall(x,y)\in S_{adj},\ \|\phi(x)-\phi(y)\|\leq d_{\text{lang}}(x,y)\implies\forall x,y\in S,\ \|\phi(x)-\phi(y)\|\leq\tilde{d}(x,y)$
@@
We follow the proof of CSD (Park et al., 2023a) similarly, but it has significant differences which will be covered in later of this section.
@@
Now we are ready to move on to the main proof. We assume that $\forall(x,y)\in S_{adj},\ \|\phi(x)-\phi(y)\|\leq d_{\text{lang}}(x,y)$. Also, note that $\forall x,y\in\mathcal{S}$, both $d_{\text{lang}}(x,x)=0$ and $d_{\text{lang}}(x,y)=d_{\text{lang}}(y,x)$ holds trivially. We denote the optimal path as $p^{*}=(x,s_{1}^{*},s_{2}^{*},...,s_{T-1}^{*},y)\in P(x,y)$ which satisfy $\tilde{d}(x,y)=D_{\text{lang}}(p^{*})$. Then,
@@
The primary distinction is that we focus exclusively on the existing path $p\in P(x,y)$ only. Here, we refer a path “exists” between state $x$ and $y$ only if state $y$ is reachable from $x$ within a finite number of transitions. During the actual optimization process, our agents encounter existing path only, so we believe this constraints align our theory more closely with the actual training scenario. Due to this specific focus, unlike from CSD, $\tilde{d}$ can no longer serves as a lower bound for $d_{\text{lang}}$.
@@
Appendix D Latent space visualization
Figure 12: Each sub-figure corresponds to a trajectory drawn from a policy conditioned on six different skills sampled randomly. In each subfigure, the left figure displays the actual paths of both the end-effector (blue) and the object (red) in the real world while the right figure illustrates the corresponding trajectory in latent space, generated using the learned $\phi$ function. A black arrow indicates the skill vector used to produce the trajectory. The numbers indicate transition step in each of the episode.
As discussed in Section 4.1, LGSD encourages the agent to spread out evenly in the latent space. The distance between two consecutive points in the latent space is regulated by the language distance, $d_{\text{lang}}$. Fig. 12 demonstrates that the orientation of the agent’s trajectory in latent space aligns well with the sampled skill vector. Since the skill vector is sampled isotropically, agent effectively spread out in the latent space. Additionally, it is noteworthy that the distance between two consecutive points in the latent space increases especially when the cube is moved. On the other hand, when the cube is stationary, there is no further extension in the latent space as can be seen in the Fig. 12 (a). This indicates that our language distance $d_{\text{lang}}$ serves as a meaningful proxy for semantic distance.
Appendix E Full algorithm of LGSD
@@
4:        Sample skill $z\sim\mathcal{N}(0,I)$
5:        while episode not terminates do
@@
8:           Query $LLM(\cdot|s,l_{\text{prompt}})$ to produce $l_{\text{desc}}(s)$ and $l_{\text{desc}}(s^{\prime})$
@@
11:           Compute embedding vector $e_{s}=f_{\text{embed}}(l_{\text{desc}}(s))$
12:           Add $\{s,a,r,s^{\prime},d_{lang}(s,s^{\prime}),e_{s},z\}$ to buffer $\mathcal{D}$
13:        end while
14:     end for
15:     for  $\{s,a,r,s^{\prime},d_{lang}(s,s^{\prime}),e_{s},z\}$ in $\mathcal{D}$ do
16:        Update $\phi$ to maximize $\mathbb{E}_{(s,z,s^{\prime})\sim\mathcal{D}}\left[(\phi(s^{\prime})-\phi(s))^{T}z+\lambda\cdot\min(\epsilon,d_{\text{lang}}(s,s^{\prime})-\lVert\phi(s)-\phi(s^{\prime})\rVert_{2}^{2})\right]$
17:        Update $\lambda$ to minimize $\mathbb{E}_{(s,z,s^{\prime})\sim\mathcal{D}}\left[\lambda\cdot\min(\epsilon,d_{\text{lang}}(s,s^{\prime})-\lVert\phi(s)-\phi(s^{\prime})\rVert_{2}^{2})\right]$
18:        Update $\pi$ using PPO with reward $r$
19:        Update $\psi$ to minimize Mean_Squared_Error between $\psi(e_{s})$ and $z$
20:     end for
21:  end for
Appendix F Implementation details
F.1 Details of downstream tasks
We used similar settings for the downstream tasks in the locomotion domain as those used in Metra (Park et al., 2023b). For AntSingleGoal, a single goal is sampled for every episode within a range of $[-50,50]$, centered at the origin. The task is considered complete when the agent reaches within a radius of 3 units around the goal, at which point the agent receives a reward of 1 and the episode terminates. For AntMultiGoal, there are four sequential goals in total. When the first goal is reached, the next goal is sampled randomly within a range of $[-7.5,7.5]$, centered at the agent’s current location. Reaching each goal gives a reward of 1.
On the other hand, we have two differences from the task used in Metra. The first difference is in the termination condition. When the agent’s torso height drops below the threshold height of 0.31 m, the episode is instantly terminated. This threshold is the default value used in the Ant environment from the official IsaacGymEnvs library. We believe this condition encourages the agent to walk rather than “roll”, which often occurs when attempting to maximize all possible state dimensions without constraints. Another difference is that we use a higher-dimensional observation space, as detailed in F.2.1. Our observation space has 39 dimensions, whereas the original Metra work has 26 dimensions, as it is based on the MuJoCo Ant environment.
We trained the low-level controller five times with different seeds for reporting final performance. We used 20,000 updated checkpoints from each training runs of low-level controller to train the high-level controller.
F.2 Observations and hyperparameters
F.2.1 Observations